If f (x)= x+6 and g (x)= x^4 what is g (f(x))
Answer by stanbon(75887)

If f (x)= x+6 and g (x)= x^4 what is g (f(x))
-------
g[f(x)] = g[x+6] = (x+6)^4 = x^4 + 4x^3*6 + 6x^2*6^2 + 4x*6^3 + 6^4
----
= x^4 + 24x^3 + 216x^2 + 864x + 1296
=============
Cheers,
Stan H.
--------------